A test in French language was conducted in a class with 15 students. The average score was 65% with a standard deviation of 20%. What is the Z-score of a student who scored 75%?

A> 0.5
B> 0.666667
C> 2.581989
D> None of the above
1
Expert's answer
2020-03-17T12:42:34-0400

"z=\frac{x-\mu}{\sigma}=\frac{75-65}{20}=0.5"

Thus, the correct choice is A